# Enabling Efficient and Flexible Interpretability of Data-driven Anomaly Detection in Industrial Processes with AcME-AD

### v1

@@ I Introduction
Anomaly detection (AD) plays a vital role in IIoT systems, enabling the timely identification of deviations from normal behavior before they escalate into costly downtime or production issues [4]. Notably, unsupervised AD techniques are particularly advantageous in this domain, as they do not necessitate pre-labeled data, often a cumbersome and resource-intensive endeavor. However, merely identifying anomalies is insufficient. Understanding the underlying causes behind them is critical for implementing effective corrective actions. This is where explainable Artificial Intelligence (XAI) comes into play [5, 6]. XAI, also known as ML interpretability, aims to elucidate the decision-making process of ML models. While some ML methods offer inherent interpretability, others, like high-performance deep neural networks, often operate as ”black boxes,” lacking transparency in their predictions. This lack of transparency can lead to hesitation in adopting ML solutions within Decision Support Systems (DSS) due to potential trust concerns [7].
@@
This paper examines how AcME-AD [8] an efficient, and model-agnostic approach to explainability that was recently developed, can be applied in industrial settings for AD. AcME-AD works well even with unsupervised AD, as it explains predictions by highlighting how changes in input features impact the resulting anomaly score. This, combined with its computational efficiency and effective result visualizations, makes AcME-AD ideal for industrial internet of things (IIoT) monitoring through decision support systems (DSS).
@@ II Related Work
Recent advancements in data and computational capabilities enabled the widespread adoption of ML within digital manufacturing. Transitioning towards Industry 5.0, the focus shifts to a human-centric paradigm, emphasizing collaboration between operators and intelligent systems [9]. Operators evolve from primarily executing tasks to leveraging their domain knowledge alongside AI-powered tools for informed decision-making. Consequently, the need for explainable and trustworthy predictions becomes paramount for building confidence in data-driven monitoring solutions, making ML interpretability a critical factor. Examples of explainable ML industrial solutions include fault detection in machinery [4], process monitoring in various industries from semiconductor [10, 11] to home appliances [12], and predictive maintenance [13].
This work focuses on interpretable anomaly detection. Prior research on this topic spans diverse areas like wireless spectrum anomalies [14], robot activity recognition [15], and ICT threat detection [16]. To balance low latency with root-cause analysis insights, we leverage AcME-AD [8] to explain predictions. This approach maintains efficiency and model-agnosticism in line with AcME [17]. However, AcME-AD is uniquely tailored for (possibly unsupervised) anomaly detection (AD) tasks, assuming only that the model provides an anomaly score as output. This combination of explainability and efficiency positions AcME-AD as an ideal solution for industrial scenarios where timely and actionable insights are essential.
@@ III Proposed Approach
The goal of this work is to prove the effectiveness of AcME-AD [8] to explain AD models in the industrial scenario. AcME-AD offers a model-agnostic approach to explainability in AD, contrasting methods tailored for specific models like Isolation Forest [18] or for Principal Component Analysis in [19]. This flexibility allows users to leverage the best-performing AD model while gaining interpretable results from AcME-AD. Furthermore, AcME-AD prioritizes computational efficiency, making it suitable for critical scenarios requiring real-time decisions. This distinguishes it from methods like SHAP [20] which can struggle in such settings. As shown in Section IV, AcME-AD delivers explanations quickly, solidifying its value for time-sensitive applications.
Consider any AD model that produces as output an anomaly score for each data point, i.e., a scalar real value indicating its level of outlyingness, and a threshold on the anomaly score, classifying scores below the threshold as normal and above as anomalous. AcME-AD produces local explanations for a data point $\mathbf{x}$ by systematically perturbing individual features, while keeping other features fixed, and observing the impact on the anomaly score. It utilizes a convex combination of four key metrics (all with range from 0 to 1) to compute $I_{j}(\mathbf{x)}$, the importance of the $j-$feature in contributing to the anomaly score assigned to $\mathbf{x}$:
@@
    Change of Predicted Class $C_{j}$: $C_{j}=1$ if feature perturbation induces a change in the classification (anomaly vs normal) of the data point $\mathbf{x}$, depending on the anomaly score and the threshold.
@@
The weights $w_{i}$ are input parameters and can be customized. Default values are set to $w_{D}=0.3,w_{C}=0.3,w_{R}=0.2,w_{Q}=0.2$. For further details we refer to [8].
AcME-AD’s what-if visualization tool (see for example Fig. 3) provides local explanation by exploring how the anomaly score would change if a single feature’s quantile value were modified while keeping others fixed. The ten most important features, ranked by decreasing importance, are displayed. The visualization features: (i) a red dashed line representing the anomaly score of the examined data point; (ii) a black solid line indicating the normal/anomalous classification threshold; (iii) small colored bubbles, indicating the anomaly scores associated with each perturbation of the feature value and (iv) large colored bubbles on the red line depicting the data point’s actual feature values. The bubble colors convey the position of the perturbed/original value within the entire dataset distribution (specifically,their quantile). By interacting with this visualization, users can understand how modifying specific feature values might impact the anomaly score and potentially transition the data point from anomalous to normal.
To identify overall important features (i.e., relevant across different anomalies), we propose the following procedure: (i) Isolate anomalies: Select all data points flagged as anomalies by the model; (ii) Localize importance: For each anomaly, employ AcME-AD to generate a local feature ranking based on importance scores; (iii) Aggregate ranks: Count how often each feature appears at each rank position across all anomalies; (iv) Visualize insights: Normalize the counts and create a stacked bar chart (see, for example Fig. 1). The x-axis represents ranking positions (1 being most important). Each bar at position k for feature j reflects the percentage of anomalies where feature j ranked kth. This allows researchers to observe which features consistently rank high across various anomalies. For many features, consider merging those ranking below 5% into an ”others” category to improve readability. This threshold can be adjusted based on specific needs.
@@ IV-A Use Case I: Chemical processes
In the first case-study, we investigate the effectiveness of AcME-AD using the Tennessee Eastman Process (TEP) dataset[21], a well-established industrial benchmark for fault and anomaly detection. This dataset, being supervised, enables us to select a high-performing AD model to explain, crucial to properly assessing our method. Since AcME-AD explains model predictions and not the underlying physical process, it is important to pick a very accurate model to leverage domain knowledge about the process for evaluation. In principle, indeed, misattributions of feature importance could be associated with shortcomings in the model rather than limitations in the interpretability approach.
TEP data are obtained by simulation using a computational model of chemical processes comprising time series originating from normal or faulty processes. Each time series contains 500 samples with 52 features. Samples are categorized into 21 classes, where Class 0 signifies normal operational states and classes 1-20 correspond to different simulated process faults. Importantly, there exists prior knowledge about the features associated with each of the first 15 faults [22], aiding the assessment of the proposed method. In our experimental setup, we consider a subset of the available dataset to reproduce an AD scenario. We randomly select 70 normal simulations and 3 faulty simulations of fault type IDV12. We select this fault type because domain knowledge about the responsible feature is available [22, 23]. Specifically, the fault IDV12 is related to the condenser cooling water input temperature, which assumes random anomalous values during the process. The true root cause feature is xmeas_11, which is the separator temperature measure. Among the trained models, Isolation Forest (IF) [24] is the best performing one, achieving an average precision score equal to 0.77. Results for other models can be found in the code repository.
@@ IV-A1 Overall feature importance
Fig. 1 displays the overall importance bar chart for TEP. Feature xmeas_11, depicted in dark blue, is correctly identified as the most or second most relevant feature in a considerable percentage of anomalies, consistest with prior knowledge. It is second in importance only to xmeas_7, which is the reactor pressure. The authors in [22], in Section 4, builded a Sign Directed Graph to model the Tennessee Eastman Process. This graph shows that reactor pressure xmeas_7 influences xmeas_11 through only another feature, suggesting that IF might leverage this causal relationship to detect anomalies.
Consistent with prior research [17, 8], KernelSHAP [20] and AcME explanations show high similarity. This is evident when comparing Fig. 1 and 2. However, computing a local explanation with AcME-AD takes $3.03$ s and with KernelSHAP $62.32$ s (as detailed in Tab. I(a)).
@@ IV-A2 Local Interpretability
Fig. 3 illustrates AcME-AD what-if tool for local explanation. For this specific anomalous prediction, the feature xmeas_11, representing the separator temperature measurement, correctly emerges as the most relevant. Its actual value, denoted by the larger green bubble, is very high with respect to the data distribution. One can notice that reducing the separator temperature would transition the point to a normal state, at least up to a certain quantile value, represented by the leftmost light blue point. After that, further decreases in the separator temperature would classify the data point as abnormal again. The only other feature that, if perturbed, would induce a change of state is xmv_9, representing the stripper steam valve, a manipulated process variable. The pattern exhibited by this feature closely resembles that of xmeas_11.
@@ IV-B Use Case II: Packaging equipment
The second industrial case-study considers the monitoring of packaging machines. The experiments are conducted on the publicly available [^2] https://zenodo.org/records/7071747 Packaging Industry Anomaly DEtection (PIADE) dataset [25], which contains unlabeled data acquired by five industrial packaging machines. We consider sequences data, where raw observations are aggregated considering 1-hour-long time windows. Because the the five machines have different operating points, in this paper, we present results obtained considering the second machine with Equipment_ID equal to 2. Results for other machines can be found in the repository. The dataset contains 2725 data points characterized by 162 features.
@@ IV-B2 Local Interpretability
As a demonstrative example, let’s consider a randomly sampled data point detected as anomalous by LODA. We generate the corresponding AcME-AD’s what-if visualization, displayed in Fig. 4. There are four features that, if perturbed, would bring the data point to a normal state. Among these, the most important is count_sum. We recall that this feature counts the number of alarms occurred within the 1-hour interval described by the data point. Its original value, represented by the larger green bubble, falls within the highest quantile of the training set. The other smaller bubbles show the anomaly scores resulting from perturbations of count_sum across various quantiles, while keeping the other features fixed. As the count_sum value decreases, depicted by the gradual change in color of the bubbles from green to blue, the data point tends toward normality. This behavior aligns with domain knowledge, where a high number of alarms typically indicates issues in the machine, thus being anomalous situations. Notably, even a slight perturbation of count_sum leads to the classification of the data point as normal. The second most relevant feature is downtime/ downtime. The anomaly score behavior closely mirrors that of count_sum. This 1-hour-sequence data point exhibits a high frequency of transitions from downtime to downtime, which is indeed atypical. The third feature which perturbation can induce a classification change is A_012. Many occurrences of this process alarm are associated with high production levels in the equipment under consideration. Consequently, frequent occurrences of this alarm can be interpreted as an indication of a normal and desirable operational situation. The current value of A_012 has a light blue color, suggesting a significantly reduced machine throughput compared to the majority of all the data points. The what-if tool points out that if the number of A_012 increases (the green bubble), the data point would return to a normal classification, which is aligned with domain expertise. The last feature that can be perturbed to normalize the sequence is idle/performance_loss, which counts the number of machine transitions from idle to a performance loss state within the interval. However, the absence of a correlation between the values assumed by this feature and the corresponding anomaly score suggests that the feature can not be effectively utilized for taking corrective actions.
@@ IV-C AcME-AD and KernelSHAP: time comparison
The de-facto standard model-agnostic method for explaining models, including unsupervised AD ones, is KernelSHAP [20]. Despite its solid theoretical foundation, the computational overhead of KernelSHAP is a recognized issue [17]. Specifically, the computational time quickly escalates with the size of the dataset used to fit the explainer, denoted as background, which is typically equal to the training set. The authors in [27] propose to reduce the size of the background by randomly sub-sampling the original dataset. However, as extensively demonstrated in [17], this operation may lead to unreliable explanations, especially in complex dataset distributions. Table I(a) reports the time [^3] For reference, experiments are conducted on AMD Ryzen 7 3700X, 3.6GHz, RAM 20GB. required to compute a local explanation of a single anomalous data point using AcME-AD and KernelSHAP on TEP. Due to limited resources, we were able to compute only the time of KernelSHAP using at most half of the training set as background. Our findings reveal that AcME-AD is an order of magnitude faster than KernelSHAP even when using the 5% as background. When utilizing the 20% and the 50% sampling, the time demand makes KernelSHAP incompatible with the integration in Decision Support Systems.
Table I(b) illustrates the elapsed time when analyzing PIADE data. In this case, the gap is less pronounced. The motivation behind this lies in the fact that, for the same number of features, the time complexity of KernelSHAP is dominated by the number of data points, whereas the complexity of AcME-AD is driven by the dimensionality of the samples (i.e., the number of features). PIADE has a high number of features (162) and a low number of samples (2725) when compared to TEP (36500). Nevertheless, once again, AcME-AD generally outperforms KernelSHAP in terms of speed, achieving comparable performance only for 25% sampling. Moreover, this discrepancy becomes substantial when exploiting the explainability method for model selection, as in Section IV-B1, where local explanations need to be generated for every predicted anomalous point.
@@
((a)) TEP dataset.
@@
((b)) PIADE dataset
TABLE I: Single explanation times for AcME-AD and KernelSHAP.

### v2

@@ I Introduction
Anomaly detection (AD) plays a vital role in IIoT systems, enabling the timely identification of deviations from normal behavior before they escalate into costly downtime or production issues [4]. Notably, unsupervised AD techniques are particularly advantageous in this domain, as they do not necessitate pre-labeled data, often a cumbersome and resource-intensive endeavor.
However, merely identifying anomalies is insufficient. Understanding the underlying causes behind them is critical for implementing effective corrective actions. This is where explainable Artificial Intelligence (XAI) comes into play [5, 6, 7]. XAI, also known as ML interpretability, aims to elucidate the decision-making process of ML models. While some ML methods offer inherent interpretability, others, like high-performance deep neural networks, often operate as ”black boxes,” lacking transparency in their predictions. This lack of transparency can lead to hesitation in adopting ML solutions within Decision Support Systems (DSS) due to potential trust concerns [8].
@@
This paper examines how AcME-AD [9] an efficient, and model-agnostic approach to explainability that was recently developed, can be applied in industrial settings for AD.
AcME-AD works well even with unsupervised AD, as it explains predictions by highlighting how changes in input features impact the resulting anomaly score. This, combined with its computational efficiency and effective result visualizations, makes AcME-AD ideal for industrial internet of things (IIoT) monitoring through decision support systems (DSS).
@@ II Related Work
Recent advancements in data and computational capabilities enabled the widespread adoption of ML within digital manufacturing. Transitioning towards Industry 5.0, the focus shifts to a human-centric paradigm, emphasizing collaboration between operators and intelligent systems [10]. Operators evolve from primarily executing tasks to leveraging their domain knowledge alongside AI-powered tools for informed decision-making. Consequently, the need for explainable and trustworthy predictions becomes paramount for building confidence in data-driven monitoring solutions, making ML interpretability a critical factor.
Examples of explainable ML industrial solutions include fault detection in machinery [4], process monitoring in various industries from semiconductor [11, 12] to home appliances [13], and predictive maintenance [14].
This work focuses on interpretable anomaly detection. Prior research on this topic spans diverse areas like wireless spectrum anomalies [15], robot activity recognition [16], and ICT threat detection [17].
To balance low latency with root-cause analysis insights, we leverage AcME-AD [9] to explain predictions. This approach maintains efficiency and model-agnosticism in line with AcME [18]. However, AcME-AD is uniquely tailored for (possibly unsupervised) anomaly detection (AD) tasks, assuming only that the model provides an anomaly score as output. This combination of explainability and efficiency positions AcME-AD as an ideal solution for industrial scenarios where timely and actionable insights are essential.
@@ III Proposed Approach
The goal of this work is to prove the effectiveness of AcME-AD [9] to explain AD models in the industrial scenario. AcME-AD offers a model-agnostic approach to explainability in AD, contrasting methods tailored for specific models like Isolation Forest [19] or for Principal Component Analysis in [20]. This flexibility allows users to leverage the best-performing AD model while gaining interpretable results from AcME-AD.
Furthermore, AcME-AD prioritizes computational efficiency, making it suitable for critical scenarios requiring real-time decisions. This distinguishes it from methods like SHAP [21] which can struggle in such settings. As shown in Section IV, AcME-AD delivers explanations quickly, solidifying its value for time-sensitive applications.
AcME-AD produces local explanations for a data point $\mathbf{x}$ by systematically perturbing individual features, while keeping other features fixed, and observing the impact on the anomaly score. It utilizes a convex combination of four key metrics (all with range from 0 to 1) to compute $I_{j}(\mathbf{x)}$, the importance of the $j-$feature in contributing to the anomaly score assigned to $\mathbf{x}$:
@@
    Change of Predicted Class $C_{j}$: $C_{j}=1$ if feature perturbation induces a change in the classification (anomaly vs normal) of the data point $\mathbf{x}$.
@@
The weights $w_{i}$ are input parameters and can be customized. Default values are set to $w_{D}=0.3,w_{C}=0.3,w_{R}=0.2,w_{Q}=0.2$.
AcME-AD’s ’what-if‘ visualization tool (see for example Fig. 3) provides local explanation by exploring how the anomaly score would change if a single feature’s quantile value were modified while keeping others fixed. The ten most important features, ranked by decreasing importance, are displayed. The visualization features: (i) a red dashed line representing the analyzed data point’s anomaly score; (ii) a black solid line indicating the normal/anomalous classification threshold, and (iii) large colored bubbles on the red line depicting the data point’s actual feature values. These bubble colors convey the position of these values within the entire dataset’s distribution (specifically, their quantile). By interacting with this visualization, users can understand how modifying specific feature values might impact the anomaly score and potentially transition the data point from anomalous to normal.
To identify overall important features (i.e., relevant across different anomalies), we propose the following procedure: (i) Isolate anomalies: Select all data points flagged as anomalies by the model; (ii) Localize importance: For each anomaly, employ AcME-AD to generate a local feature ranking based on importance scores; (iii) Aggregate ranks: Count how often each feature appears at each rank position across all anomalies; (iv) Visualize insights: Normalize the counts and create a stacked bar chart (see, for example Fig. 1).
The x-axis represents ranking positions (1 being most important). Each bar at position k for feature j reflects the percentage of anomalies where feature j ranked kth. This allows researchers to observe which features consistently rank high across various anomalies. For many features, consider merging those ranking below 5% into an ”others” category to improve readability. This threshold can be adjusted based on specific needs.
@@ IV-A Use Case I: Chemical processes
In the first case-study, we investigate the effectiveness of AcME-AD using the Tennessee Eastman Process (TEP) dataset[22], a well-established industrial benchmark for fault and anomaly detection. This dataset, being supervised, enables us to select a high-performing AD model to explain, crucial to properly assessing our method. Since AcME-AD explains model predictions and not the underlying physical process, it is important to pick a very accurate model to leverage domain knowledge about the process in evaluating interpretability. In principle, indeed, misattributions of feature importance could be associated with shortcomings in the model rather than limitations in the interpretability approach.
TEP data are obtained by simulation using a computational model of chemical processes. The dataset comprises time series originating from normal processes or faulty processes. Each time series contains 500 samples with 52 features. Samples are categorized into 21 classes, where Class 0 signifies normal operational states and classes 1-20 correspond to different simulated process faults. Importantly, there exists prior knowledge about the features associated with each of the first 15 faults [23], aiding the assessment of the proposed method.
In our experimental setup, we consider a subset of the available dataset to reproduce an AD scenario. We randomly select 70 normal simulations and 3 faulty simulations of fault type IDV12. We select this fault type because domain knowledge about the responsible feature is available [23, 24]. Specifically, the fault IDV12 is related to the condenser cooling water input temperature, which assumes random anomalous values during the process. The true root cause feature is xmeas_11, which is the separator temperature measure. Among the trained models, Isolation Forest (IF) [25] is the best performing one, achieving an average precision score equal to 0.77. Results for other models can be found in the code repository.
@@ IV-A1 Overall feature importance
Fig. 1 displays the overall importance bar chart for TEP. Feature xmeas_11, depicted in dark blue, is correctly identified as the most or second most relevant feature in a considerable percentage of anomalies, consistest with prior knowledge. It is second in importance only to xmeas_7, which is the reactor pressure. The authors of [23], in Section 4, builded a Sign Directed Graph to model the Tennessee Eastman Process. This graph shows that reactor pressure xmeas_7 influences xmeas_11 through only another feature, suggesting that IF might leverage this causal relationship to detect anomalies.
Consistent with prior research [18, 9], KernelSHAP [21] and AcME explanations show high similarity. This is evident when comparing Figures 1 and 2. However, computing a local explanation takes $3.03$ s with AcME-AD and $62.32$ s with KernelSHAP (as detailed in Tab. I).
@@ IV-A2 Local Interpretability
Fig. 3 illustrates AcME-AD what-if tool for local explanation. For this specific anomalous prediction, the feature xmeas_11, representing the separator temperature measurement, correctly emerges as the most relevant. Its actual value, denoted by the larger green bubble, is very high. However, one can notice that reducing the separator temperature would transition the point to a normal state, at least up to a certain quantile value, represented by the leftmost light blue point. After that, further decreases in the separator temperature would classify the data point as abnormal again. The only other feature that, if perturbed, would induce a change of state is xmv_9, representing the stripper steam valve, a manipulated process variable. The pattern exhibited by this feature closely resembles that of xmeas_11.
@@ IV-B Use Case II: Packaging equipment
The second industrial case-study considers the monitoring of packaging machines. The experiments are conducted on the publicly available [^2] https://zenodo.org/records/7071747 Packaging Industry Anomaly DEtection (PIADE) dataset [26], which contains unlabeled data acquired by five industrial packaging machines. We consider sequences data, where raw observations are aggregated considering 1-hour-long time windows. Because the the five machines have different operating points, in this paper, we present results obtained considering the second machine with Equipment_ID equal to 2. Results for other machines can be found in the repository. The dataset contains 2725 data points, each one characterized by 162 features.
@@ IV-B2 Local Interpretability
As a demonstrative example, let’s consider a randomly sampled data point detected as anomalous by LODA. We generate the corresponding AcME-AD’s what-if visualization, displayed in Fig. 4. There are four features that, if perturbed, would bring the data point to a normal state. Among these, the most important is count_sum. We recall that this feature counts the number of alarms occurred within the 1-hour interval described by the data point. Its original value, represented by the larger green bubble, falls within the highest quantile of the training set, being anomalous on its own. The other smaller bubbles show the anomaly scores resulting from perturbations of count_sum across various quantiles, while keeping the other features fixed. As the count_sum value decreases, depicted by the gradual change in color of the bubbles from green to blue, the data point tends toward normality. This behavior aligns with domain knowledge, where a high number of alarms typically indicates issues in the machine, thus being anomalous situations. Notably, even a slight perturbation of count_sum leads to the classification of the data point as normal.
The second most relevant feature is downtime/ downtime. The anomaly score behavior closely mirrors that of count_sum. This 1-hour-sequence data point exhibits a high frequency of transitions from downtime to downtime, which is indeed atypical.
The third feature which perturbation can induce a classification change is A_012. Many occurrences of this process alarm are associated with high production levels in the equipment under consideration. Consequently, frequent occurrences of this alarm can be interpreted as an indication of a normal and desirable operational situation. The current value of A_012 has a light blue color, suggesting a significantly reduced machine throughput compared to the majority of all the data points. The what-if tool points out that if the number of A_012 increases (the green bubble), the data point would return to a normal classification, which is aligned with domain expertise.
The last feature that can be perturbed to normalize the sequence is idle/performance_loss, which counts the number of machine transitions from idle to a performance loss state within the interval. However, the absence of a correlation between the values assumed by this feature and the corresponding anomaly score suggests that the feature can not be effectively utilized for taking corrective actions.
@@ IV-C AcME-AD and KernelSHAP: time comparison
The de-facto standard model-agnostic method for explaining models, including unsupervised AD ones, is KernelSHAP [21]. Despite its solid theoretical foundation, the computational overhead of KernelSHAP is a recognized issue [18]. Specifically, the computational time quickly escalates with the size of the dataset used to fit the explainer, denoted as background, which is typically equal to the training set. The authors of [28] propose to reduce the size of the background by randomly sub-sampling the original dataset. However, as extensively demonstrated in [18], this operation may lead to unreliable explanations, especially in complex dataset distributions. Table I reports the time [^3] For reference, experiments are conducted on AMD Ryzen 7 3700X, 3.6GHz, RAM 20GB. required to compute a local explanation of a single anomalous data point using AcME-AD and KernelSHAP on TEP. Due to limited resources, we were able to compute only the time of KernelSHAP using at most half of the training set as background. Our findings reveal that AcME-AD is an order of magnitude faster than KernelSHAP even when using the 5% as background. When utilizing the 20% and the 50% sampling, the time demand makes KernelSHAP incompatible with the integration in Decision Support Systems.
Table II illustrates the elapsed time when analyzing PIADE data. In this case, the gap is less pronounced. The motivation behind this lies in the fact that, for the same number of features, the time complexity of KernelSHAP is dominated by the number of data points, whereas the complexity of AcME-AD is driven by the dimensionality of the samples (i.e., the number of features). PIADE has a high number of features (162) and a low number of samples (2725) when compared to TEP (36500). Nevertheless, once again, AcME-AD generally outperforms KernelSHAP in terms of speed, achieving comparable performance only for 25% sampling. Moreover, this discrepancy becomes substantial when exploiting the explainability method for model selection, as in Section IV-B1, where local explanations need to be generated for every predicted anomalous point.
@@
TABLE I: Single explanation times for AcME-AD and KernelSHAP with TEP dataset.
@@
TABLE II: Single explanation times for AcME-AD and KernelSHAP with PIADE dataset.